OpenStudy (anonymous):

Which equation represents a graph with a vertex at (1, –6)?

OpenStudy (anonymous):

Which equation represents a graph with a vertex at (1, –6)? y = 3x2 + 6x – 3 y = 3x2 – 6x – 3 y = 3x2 – 8x – 1 y = 3x2 – 3x –

Nnesha (nnesha):